QUESTION 4
The revenue function of a company is R = 36q – 2q², where p represent the price per unit of the
good and q represent the quantity demanded. Find the maximum value of the revenue for the revenue function given.
QUESTION 5
Find the value of x for the following logarithm functions:
a) log2 16^x + log2 8 = -½
b) log3 x^2 + log3 4 = log3 1

Answer by stanbon(48535)

QUESTION 4
The revenue function of a company is R = 36q – 2q², where p represent the price per unit of the good and q represent the quantity demanded. Find the maximum value of the revenue for the revenue function given.
--------
Max occurs where q = -b/2a = -36/2*-2 = 9
R(9) = 36*9-2*9^2 = 162
---------------------------
QUESTION 5
Find the value of x for the following logarithm functions:
a) log2 16^x + log2 8 = -½
log2 2^4x + log3 2^3 =
4x+3 = -1/2
8x+6 = = -1
8x = -7
x = -7/8
============
b) log3 x^2 + log3 4 = log3 1
log2 [4x^2] = log3 1
4x^2 = 1
x^2 = 1/4
x = 1/2 or x = -1/2
==================
Cheers,
Stan H.